# Learning Prehensile Dexterity by Imitating and Emulating State-only Observations

### v1

@@ 16      Update the Motion Refinement Policy $\Psi$ with $\mathcal{B}$;
Appendix D Details of KODex Training
D-A Modeling Dexterous Manipulation Skills
A central principle behind KODex is that the desired behavior of a robot can be represented using a dynamical system. To this end, we define the state at time $t$ as $\mathrm{x}(t)=[{\mathrm{h}_{t}}^{\top},{\mathrm{o}_{t}}^{\top}]^{\top}$, where ${\mathrm{h}_{t}}\in\mathcal{H}\subseteq\mathbb{R}^{n}$ and ${\mathrm{o}_{t}}\in\mathcal{O}\subseteq\mathbb{R}^{m}$ represent the state of the robot and the object, respectively, at time $t$. As such, the dynamical system we wish to capture is
@@ D-A Modeling Dexterous Manipulation Skills
D-B Learning the Koopman Matrix $\mathbf{K}$
The next step is to learn the Koopman matrix $\mathbf{K}$ from dataset $\mathcal{D}=[\{{\mathrm{h}}_{t}^{(1)},{\mathrm{o}}_{t}^{(1)}\}^{t=T^{(1)}}_{t=1},\cdots,\{{\mathrm{h}}_{t}^{(N)},{\mathrm{o}}_{t}^{(N)}\}^{t=T^{(N)}}_{t=1}]$ (Section. III-B). Recall that CIMER’s Motion Generation Policy $\Phi$ is formulated as the linear dynamical system shown in (3).
@@ D-B Learning the Koopman Matrix $\mathbf{K}$
$\mathbf{A}=\sum_{n=1}^{n=N}\sum_{t=1}^{t=T^{(n)}-1}\frac{g(\mathrm{x}^{(n)}(t+1))\otimes g(\mathrm{x}^{(n)}(t))}{N(T^{(n)}-1)},\ \mathbf{G}=\sum_{n=1}^{n=N}\sum_{t=1}^{t=T^{(n)}-1}\frac{g(\mathrm{x}^{(n)}(t))\otimes g(\mathrm{x}^{(n)}(t))}{N(T^{(n)}-1)}$ (4)
@@
Appendix E Emulation might not be necessary for Non-prehensile Manipulation
@@
Appendix F CIMER Hyperparameters
TABLE II: Training details
[…]
$(256,128)$ 2e-6 0.98
[…]
0.2 8 0.97
In addition, for all tasks, we used $k^{h}=5,k^{o}=5,r_{t}^{b}=25$ to compute the tracking rewards described in Sec. III-D.

### v3

@@ 16      Update the Motion Refinement Policy $\Psi$ with $\mathcal{B}$;
Appendix E Details of KODex Training
Modeling Dexterous Manipulation Skills: A central principle behind KODex is that the desired behavior of a robot can be represented using a dynamical system. To this end, we define the state at time $t$ as $\mathrm{x}(t)=[{\mathrm{h}_{t}}^{\top},{\mathrm{o}_{t}}^{\top}]^{\top}$, where ${\mathrm{h}_{t}}\in\mathcal{H}\subseteq\mathbb{R}^{n}$ and ${\mathrm{o}_{t}}\in\mathcal{O}\subseteq\mathbb{R}^{m}$ represent the state of the robot and the object, respectively, at time $t$. As such, the dynamical system we wish to capture is
@@ Appendix E Details of KODex Training
Learning the Koopman Matrix $\mathbf{K}$: The next step is to learn the Koopman matrix $\mathbf{K}$ from dataset $\mathcal{D}=[\{{\mathrm{h}}_{t}^{(1)},{\mathrm{o}}_{t}^{(1)}\}^{t=T^{(1)}}_{t=1},\cdots,\{{\mathrm{h}}_{t}^{(N)},{\mathrm{o}}_{t}^{(N)}\}^{t=T^{(N)}}_{t=1}]$ (Section. III-B). Recall that CIMER’s Motion Generation Policy $\Phi$ is formulated as the linear dynamical system shown in (3).
@@
$\begin{split}\mathbf{A}&=\sum_{n=1}^{n=N}\sum_{t=1}^{t=T^{(n)}-1}\frac{g(\mathrm{x}^{(n)}(t+1))\otimes g(\mathrm{x}^{(n)}(t))}{N(T^{(n)}-1)},\\ \mathbf{G}&=\sum_{n=1}^{n=N}\sum_{t=1}^{t=T^{(n)}-1}\frac{g(\mathrm{x}^{(n)}(t))\otimes g(\mathrm{x}^{(n)}(t))}{N(T^{(n)}-1)}\end{split}$ (4)
@@
Appendix F Emulation might not be necessary for Non-prehensile Manipulation